5. For this problem you will analyze the elasticity of substitution and the isoquant graphs for the following production func(c) What is the elasticity of substitution for F? (3 points)
Answer #1

Q = 2K1/2 L1/2

Put Q = 10

10 = 2K1/2 L1/2

(10 / 2)= K1/2 L1/2

5 = K1/2 L1/2

(5 /  L1/2 ) = K1/2

(5 /  L1/2 )2 = (K1/2)2

(25 / L) = K (Isoquant equation for output level of 10)

(25 /L ) =K
L K
1 25.00
2 12.50
3 8.33
4 6.25
5 5.00
6 4.17
7 3.57
8 3.13
9 2.78
10 2.50
11 2.27
12 2.08
13 1.92
14 1.79
15 1.67
16 1.56
17 1.47
18 1.39
19 1.32
20 1.25

20.00 19.00 18.00 Capital 900 N 5. K=5 on 1 =10. K-2.5 0.00 0 1 2 3 4 5 6 7 8 11 12 13 14 15 16 17 18 19 20 9 10 Labor

(b) Marginal Rate of Technical Substitution (MRTS) = (MPL / MPK)

MP L = \Delta F(K,L)/ \Delta L

MPL = 2 * (1/2) K1/2 L-1/2

MPL = K1/2 L-1/2

MP K = \Delta F(K,L)/ \Delta K

MPK= 2 * (1/2) K-1/2 L1/2

MPK= K-1/2 L1/2

MRTS = MPL /MPK

MRTS = ( K1/2 L-1/2 / K-1/2 L1/2 )

MRTS = K / L

(c)

Elasticity of substitution (o): o= -O(K/L) MRTS OMRTS (K/L)

we have, MRTS = K/L

\Delta MRTS / \Delta (K/L) = 1

\Delta (K/L)/\Delta MRTS = 1

Elasticity of substitution = 1 * (MRTS / MRTS) (Note: MRTS = K/L)

Elasticity of substitution = 1